Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
mantovani, valentini, vardaro - 19:20 Thursday 02 December 2021 (54082) Print this report
ISC shift report: carm offset zero and SR alignment study

Today' shift focused on checking the reproducibility of yesterday's long lock (54067) and to further check the influence of SR misalignment on the SDB1 and SDB2 quadrant signals.

Three locks at CARM offset zero were achieved:

  • 9.15-9.40 UTC, ~25 minutes long, a small tuning of SR TY (0.6 rad) was performed and triggered a large improvement in the dark fringe stability and level at 9.18 utc (see figure 1 and 2). A sudden ~85 Hz oscillation noticeable on DARM and DIFFp TY led to the unlock (figure 3).
  • 11.26-11.38 : the engagement of SSFS and the 1f handoffs of PRCL and MICH were tested successfully during this lock.  In the second part of the lock the alignment of SR TX was moved in steps of 0.2 rad each (see figure 4).
  • 12.45-12.58 : further misalignment steps of SR were tested in this lock, both in TX and TY (see figure 5 and 6)

Further attempts failed due to difficulties in engaging the lock of the CITF in the afternoon.

 

Some notes on the SR misalignment steps and CARM offset zero behaviour:

  • In today's locks, no clear effect of the SR misalignment was noticed neither on the SDB1 LC (closed on the SDB1_B5_QD2 DC quadrant) nor on the SDB2 56MHz quadrant signals (see figure 2-6). This is in opposition to what observed in yesterday's long lock (see figure 7)  and in the 30th november locks (54061 ). The reason of this change of behaviour is still under investigation.
  • In the first lock, the SR alignment change 'triggered' a  (lately recurrent) sudden change in the ITF behaviour. This effect has been observed several times in the past days too, and often happens a couple of minutes after reaching carm offset zero without any particular operation triggering it (e.g. today at 12.46.20 UTC, see figure 8).  The usual features of this behaviour change on many signals, including the following:
    • on B1p, B4, B7, B8 DC signals
    • on the B1p and B5 galvo and quadrant signals,
    • beam shape observed on cameras and phase cameras
    • DARM error signal and correction spectra (this also leads to an offset change on the test mass AA centering signals obtained from demodulating the dithering lines from the DARM correction sent to the input mirrors)
    • DIFFp error signals
    • a reduction of slope of the B4_12 and B4_112 MHz signals
    • sidebands balance

Additional operations performed:

  • The engagement of the SDB1 drift control using the LC_B5_QD2_enbl relay has been added in the automation in DRMI_LOCK.py. It is now automatically engaged once the 3f handoff of the CITF is performed, and it is disabled in the DOWN state.
  • The demodulation phase of B2_169MHz (used for MICH and SRCL) has been reduced by ~0.5 rad in the 160mW and 480mW locking steps, in order to avoid unlocks during the CARM offset reduction procedure.

 

The ITF was left in DOWN state in order to proceed with the TCS calibration activities.

 

 

Images attached to this report
Comments to this report:
valentin - 20:45 Thursday 02 December 2021 (54084) Print this report

An additional note:

As noticed by Bonnand (in 54075) and Ruggi, yesterday's arm transmitted power at CARM offset zero was lower than usual (actually it was the lowest since a few weeks), most likely due to a misalignment of the test masses (see figure). In today's attempts, the cavity power was higher, together with the sideband decay rate. Figure 1 shows a comparison between yesterday and today's locks, including the (currently out of loop) test-mass centering dithering error signals.

The possible (positive) effect of yesterday's lower arm circulating power (due to the misalignment) on the sideband decay rate and lock longevity should be further evaluated.

Images attached to this comment
mwas - 8:37 Friday 03 December 2021 (54089) Print this report

The effects of SR TY alignment are still visible on B5 QD2 56MHz signal during the Dec 2 locks, but less clearly. One needs to look at the trend data to remove some of the high frequency noise.

Figure 1 shows the lock around 9:15 UTC.  The 0.4 urad step on SR TY makes a clear step on B5_QD2_H_56MHz, and also on SDB1 TX that follows the B5 DC beam. The latter is surprising as this is the vertical direction while we move SR horizontally.

Figure 2 shows the lock around 11.25 UTC. The 0.2 urad steps on SR TX do not show any clear signal on the B5 quadrant signals.

Figure shows the lock around 12.45:UTC. The series of 0.2urad steps on SR TY seems to increase the error on B5 QD2 H 56MHz this is less clear than in figure 1, but that could be because we are moving away from the zero of the signal. The starting point was already at 0.02V on the quadrant, which in Nov 30 data corresponded to the end of the linear range of the quadrant error signal sensitivity to SR (and to a ~1urad misalignment).

In conclusion B5_QD2_H_56MHz still seems to be a good candidate for SR TY control. For SR TX the vertical signal of the same quadrant doesn't seem to have visible information.

Images attached to this comment
rocchi - 15:23 Friday 03 December 2021 (54091) Print this report

I gave a closer look at the data reported in this entry. In particular to the first 10 minutes since the lock is acquired, because there is where a lot of the thermal lensing occurs (see figure 1): after 10 minutes, the curvature is already 55% of the steady state.

Checking the values of B7 for the four locks under investigation (again after 10 minutes), the intracavity power during the December 1st long lock is lower with respect to the other locks by less than 2% (figure 2).

Furthermore, I made a fit of the B4_112MHz_mag data (figure 3) with an exponential function to extract the decay time during these four locks. Figure 4 shows an example of such fits, where the outliers have been removed and only the decading transient has been considered. The results are reported in the table below.

Lock date Decay time [s]
21-12-01 13:04 UTC 2.8e3
21-12-02 09:14 UTC 4.3e3
21-12-02 11:26 UTC 1.7e3
21-12-02 12:44 UTC 2.6e3

So, the conclusion is that, at least during the transient (first 10 minutes of the locks), there is no such big difference in the intracavity power during these locks, and that the decay times of the 56 MHz sidebands are not correlated with the power in the arms.

Images attached to this comment
Search Help
×

Warning

×